hello, My name is Jasmine
I am trying to solve with this equation c(g)=3.03(g).
what does this number represent 3.03
find c(2)
find c (9)
how do I work this

Answer by Alan3354(69443)

I am trying to solve with this equation c(g)=3.03(g).
what does this number represent 3.03
find c(2)
find c (9)
how do I work this
---------------
I think it's c(g) = 3.03*g (3.03 times g)
3.03 is a coefficient, for a multiplier.
It's also the slope of the graph.
-----------------
For c(2), sub 2 for g
c(2) = 3.03*2 = 6.06
c(9) = 3.03*9 = 27.27